# Out-of-Distribution Detection: A Task-Oriented Survey of Recent Advances

### v3

@@ 3.2. OOD Detection Approaches with Both ID and OOD Data
Outlier Mining. The traditional OE concept assumes the existence of ID input $\mathcal{D}_{\text{in}}$ and OOD input $\mathcal{D}_{\text{out}}$, both independently and heterogeneously distributed, originating from different sources. However, this premise cannot be fully guaranteed in the current training process due to potential noise in the training OOD data. Outlier Mining differs slightly from the traditional OE approach in that, although it also utilizes real-world OOD samples to address the issue, it focuses on identifying the optimal selection within the existing OOD data. The main process is depicted in Fig.  3 (d).
POEM  (Ming et al., 2022a) raises methodology leans towards excavating anomalies that carry more representative characteristics.  Li and Vasconcelos (2020) propose a method of data resampling to acquire representative outlier data for training, assigning higher priority score reweighting to hard negative instances. Iterative optimization, leveraging adversarial principles, selects target data. POEM employs posterior sampling to unearth anomalies with elevated boundary scores from an extensive auxiliary dataset, facilitating a nuanced comprehension of the intricacies within OOD samples. DAOL  (Wang et al., 2023a) takes the disparity between authentic data and auxiliary OOD data as a starting point to formulate an OOD dataset. Utilizing the Wasserstein ball, it models the distribution of all OOD data, selecting the most challenging OOD data within the ball for training.
@@
Long-tail ID data. Given practical requirements, research is progressively increasing regarding scenarios where ID data is imbalanced during training yet still provides OOD data information. The concept behind PASCAL  (Wang et al., 2022b) is to segregate tail-end data and OOD data solely through contrastive loss during training phase, aiding the model in better distinguishing between the two. To address the issue of model confusion between OOD samples and tail-end class data, COCL  (Miao et al., 2024a) incorporates a learnable tail class prototype during the training process. This prototype serves to bring tail-end samples closer while distancing them from OOD data, thereby mitigating the model’s bias towards OOD samples.  Choi et al. (2023) take a different approach, suggesting that factors affecting the performance of OOD detection may be related to the imbalance in the cross-class distribution of auxiliary OOD data. Consequently, it proposes an energy regularization loss, specifically regularizing auxiliary samples from the majority classes to address the issue of class imbalance in OOD data. EAT  (Wei et al., 2024) utilizes dynamically assigned virtual labels during the model training process to train OOD data, thereby expanding the classification space.
@@ 4.1. Post-hoc OOD Detection Approaches
Overview. Given a well-trained model, this problem scenario involves utilizing only the intermediate results computed by the trained model during testing, without modifying any parameters of the model, to accomplish the OOD detection task. Post-hoc methods are favored for their lightweight nature, low computational costs, and the fact that they require minimal modifications to the model and objectives. Its main objective is to construct an effective scoring function that can accurately reflect the behavior of ID data. These characteristics make them highly desirable for convenient and straightforward deployment in practical scenarios.
Post-hoc approaches are categorized into five types: Output-based, Distance-based, Gradient-based, Feature-based and Density-based. Recent work on this type of problem has some recent progress. A summary of the key factors involved in such methods is given in Table  2.
@@
Output-based. Algorithms based on output primarily aim to explore the latent representations of the output from the intermediate layers of neural networks, which include logits and class distributions, among others. MSP  (Hendrycks and Gimpel, 2017) is the first to employ the maximum softmax value to validate OOD detection effectiveness. For OOD samples, their output probability distribution tends to be closer to a uniform distribution, demonstrating the model’s inability to correctly classify the category. Departing from the MSP method, MaxLogits  (Hendrycks et al., 2022) detects anomalies by utilizing the maximum value of the logits vector generated by the neural network. These logits constitute the raw, pre-Softmax layer outputs of the neural network, representing the model’s confidence in each category without undergoing Softmax transformation. Alternatively, the Energy  (Liu et al., 2020a) method defines an energy score conceptually derived from Helmholtz Free Energy. This score can theoretically reflect the input probability density. By assigning lower energy scores to ID data, this method distinguishes them from anomalous data. Additionally, it demonstrates reduced susceptibility to model overconfidence issues.
@@
Gradient-based. Grad  (Lee and AlRegib, 2020) research indicates that gradient-based methods also contribute to OOD detection by quantifying model uncertainty through the gradients propagated in reverse during backpropagation. If the input sample is an ID sample, the gradients of the model with respect to these samples tend to be relatively small and stable. Conversely, the gradients are typically larger or more irregular for OOD samples. GradNorm  (Huang et al., 2021) posits that the gradient magnitude for ID data surpasses that of OOD data. Leveraging this observation, it employs gradient vector norms, computing KL divergence through softmax output and uniform distribution backpropagation to detect OOD data. In contrast, GradOrth  (Behpour et al., 2023) adopts an alternative perspective, recognizing that crucial features of OOD data reside in a low-rank subspace. Consequently, it shifts focus to calculating gradient projection norms in this subspace to identify OOD data. GAIA  (Chen et al., 2023) employs a combination of gradient anomaly checks and aggregation. It enables the model to interpret indeterminacy attributively, introducing channel-wise average abnormality and zero-deflation abnormality without prior knowledge to gauge the extent of data distribution variations. The OPNP  (Chen et al., 2024a) method discovers that the OOD detection capability of the model is highly sensitive to parameters and activated neurons that are close to zero. Therefore, this method utilizes a pruning behavior for parameters and neurons to remove those leading to overfitting, thereby enhancing the model’s generalization ability. In recent research, PRO (Chen et al., 2025a) introduces an adversarial score, with the core idea being to perturb the input data using gradient descent to search for the local minimum score near the original input. This method effectively lowers the confidence of OOD samples and is compatible with existing MSP detection methods during score calculation, thereby enhancing detection performance. S&I (Zhang et al., 2025a) identifies the issue of insufficient sample gradient feature discrimination in this type of method. By introducing adversarial samples and integrating gradients along the attribution path, it achieves a more accurate feature distinction strategy.
Feature-based. Another avenue of research into OOD detection involves the impact of neural network intermediate variables relative to the final prediction. Inspired by adversarial examples, ODIN  (Liang et al., 2018a) experimentally discovered that adding small perturbations to the input image features can more effectively detect OOD data. The perturbation formula is as follows:
@@
where the parameter $\varepsilon$ is the perturbation magnitude. For a given input $x$, compute its logit output $p_{c}(x)$. Therefore, ReAct  (Sun et al., 2021) focuses on the high activation values in the intermediate results of the model. These activation values do not affect the model’s classification, but truncating high activation values can significantly improve OOD detection performance. VRA  (Xu et al., 2023), an extended iteration of ReAct  (Sun et al., 2021), builds upon the premise of ReAct which truncates only high activation values. However, VRA posits that this might not be the optimal solution and, therefore, employs a variational approach to seek an optimal solution. It utilizes piecewise functions to emulate suppression or amplification operations, aiding the model in recognizing anomalous data. KANs (Canevaro et al., 2025) also focus on the neuron activation states. Due to the local neuroplasticity characteristic of Kolmogorov-Arnold Networks (Liu et al., 2025b), KANs distinguish between ID and OOD data by comparing the activation response differences between the trained and untrained KANs. SHE  (Zhang et al., 2022) converts the output of the penultimate layer into a storage pattern, utilizing Hopfield energy  (Krotov and Hopfield, 2016) for OOD sample detection through a pattern of storing before comparison. It employs a storage mechanism to calculate the average logits for each class, preparing this simple storage pattern for subsequent OOD detection tasks. Drawing inspiration from experience, DDCS  (Yuan et al., 2024) adapts to select suitable channels for data classification after correcting each channel in the neural network. These channels are evaluated based on inter-class similarity and variance to measure their discriminative power for ID data. LINe  (Yong Hyun Ahn, 2023) similarly emphasizes neuron outputs at the feature level. It uses Shapley value pruning to select only high-contribution neurons for prediction while masking out the remaining input data, thereby reducing the influence of irrelevant outputs. CADRef (Ling et al., 2025) decouples the features by using a strategy based on the sign alignment between the relative features and model weights, dividing the sample features into positive and negative errors. Intuitively, the positive and negative error components influence the final output logits, which helps in identifying OOD images from this perspective. ITP (Xu and Yang, 2025) optimizes the model by quantifying the contribution of model parameters to ID data prediction, specifically by evaluating the contribution of each parameter through its partial derivatives and removing those with low contribution. Additionally, during the testing phase, ITP uses a right-tail Z-score test to assess whether any parameter exhibits overconfidence in the classification of a given test sample, further refining the model’s decision-making.
Feature shaping is emerging as a prominent area of research focus. This methodology entails the refinement of intermediate representations—specifically, the intermediate features—within the forward propagation phase of the model. The salient advantage of this approach lies in its non-disruptive nature concerning the original classification results, coupled with its simplicity and efficacy. Building upon this foundation, ViM  (Wang et al., 2022c) simultaneously considers the roles of features, logits, and probabilities, constructing virtual logits to aggregate information for decision-making. By decomposing the features of the penultimate layer in a neural network, it identifies the null space that is irrelevant to classification but exhibits exceptional performance in OOD detection. Its computation of this score $S$ can be expressed in the following form:
@@
where $\alpha$ is a scaling constant, computed by the model. Here $z=z^{P}+z^{P^{\perp}}$ and $z^{P^{\perp}}$ is the projection of $z$ to $P^{\perp}$. And it have $Wz^{P^{\perp}}=0$. Additionally, $LogSumExp$ represents the computation process of the energy function  (Liu et al., 2020a), and $f(z)$ represents the logit output of the model. The first term here represents virtual logits, while the second term represents the score of the energy function. Neco  (Ammar et al., 2024) subsequently reveals the prevalent phenomenon of neural collapse in contemporary neural networks, impacting OOD detection performance. The observation of orthogonal trends between ID data and OOD data features is leveraged to differentiate OOD data. Similarly, NCI (Liu and Qin, 2023) studies the phenomenon of neural collapse and discovers that during training, ID samples tend to cluster around class weight vectors in the penultimate feature space, while OOD samples do not exhibit this clustering behavior. Therefore, NCI leverages the proximity between features and class weight vectors to effectively distinguish ID samples from OOD samples, while maintaining low computational cost. ASH  (Djurisic et al., 2023) is a straightforward method for dynamic activation modification where a substantial portion of activations in samples during later stages of model training is either removed or slightly adjusted. NAC (Liu et al., 2024b) introduces a measure of neuron activation coverage, based on the premise that if a neuron in a neural network is rarely activated, this state may indicate a higher likelihood of OOD data. By quantifying this statistical property, NAC aims to distinguish between ID and OOD data. Based on the existing literature on feature shaping,  Zhao et al. (2024) utilize a piecewise constant shaping function to partition the feature domain into disjoint intervals, estimating a scalar within each interval as an approximation. As the interval width approaches zero, an approximation of the maximum logits can be obtained. The goal of BLOOD  (Jelenić et al., 2024) is to smooth the representations of intermediate layers in neural networks for predicting OOD data. This study found that, compared to OOD data, ID data exhibits smoother variations in the intermediate layer representations of neural networks. Leveraging this characteristic, new statistical measures can be designed to discriminate against anomalous data. In addition, another line of work focused on neuron activation pruning, which is a new activation shaping scheme proposed based on the research foundation of ASH. SCALE  (Xu et al., 2024a) emphasizes scaling as a crucial metric for evaluating samples, and it similarly finds significantly lower pruning rates for OOD data. Therefore, the proposed reshaping of intermediate layer tensors can effectively enhance detection performance.
Density-based. Recent advancements in density-based OOD detection models have demonstrated substantial performance gains, predicated on the models’ ability to accurately capture and understand the intrinsic characteristics of the true data distribution. For example, GEM  (Morteza and Li, 2022) models the feature space of ID data as a class-conditional multivariate Gaussian distribution. Under this assumption, it designs new statistical metrics to validate the model’s performance. Using a modeling approach based on Gaussian mixture models, GEM Score is aligned with the true log-likelihood for capturing OOD uncertainty. However, while GEM strictly relies on Gaussian assumptions, recent works ConjNorm  (Peng et al., 2024) introduce a novel framework based on Bregman divergence, extending the consideration of data distributions to encompass an exponential family of distributions. By redefining density functions for data, this model has a broader range of applications.
@@ 4.2. Test-Time Adaptive OOD Detection Approaches
Overview. Test-time adaptive methods, which employ a classifier trained on the training set, strive to utilize unlabeled test data, either the complete test set or a series of mini-batches, to enhance OOD detection performance through model adaptation. Test-time adaptive approaches are based on the theoretical insight  (Fang et al., 2022) that detecting OOD samples using only ID samples without any additional knowledge can be challenging and potentially limited. These methods can be divided into two categories according to whether the model be modified during the testing time: Model-optimization-based and Model-optimization-free. Both of them undergo a post-training phase, during which the trained model can be adapted, regardless of whether it is updated.
@@
A line of these methods (Katz-Samuels et al., 2022; Du et al., 2024a) advocate for harnessing a combination of unlabeled ID and OOD data, termed ”wild data”, which is plentiful and readily available in the real-world scenario, to OOD detection. The motivation of WOODS  (Katz-Samuels et al., 2022) is to clean the wild data to get reliable OOD candidates, and then the model regularization can be performed with the knowledge of them. Afterwards, to understand the role of wild data in OOD detection, SAL  (Du et al., 2024a) explains how they help OOD detection from the lens of separability and learnability. Notably, wild data is not entirely equivalent to test OOD data if wild OOD data is not from the test datasets, inevitably leading to incorporating unintended information into the model. In addition, the extra training requirement in WOODS, thought necessary, incurs significant costs and is generally undesirable.
Another line of model-optimization-based approaches draws inspiration from the semi-supervised-learning (SSL) techniques  (Zhu, 2005), aiming for a more efficient, lightweight training process during the post-training phase. Pseudo-labeling  (Arazo et al., 2020) is a simple but effective way to label the test data, which enhances learning from test unlabeled data. The method proposed by  Yang et al. (2023a), termed AUTO, employs only pseudo-OOD data to refine the model. The role of pseudo-ID data in AUTO is to mitigate catastrophic forgetting with a semantically-consistent objective, thereby maintaining the accuracy of ID classification. In contrast, ATTA  (Gao et al., 2023a) and SODA  (Geng et al., 2023) harness both pseudo-ID and pseudo-OOD data to refine the trained model. SODA employs a dual-loss approach to tackle pseudo-ID and pseudo-OOD data simultaneously, while ATTA distinguishes them with different weighting techniques.
@@
Model-optimization-free. Modifying the original trained model is infeasible in certain security-sensitive scenarios. Therefore, methods enabling test-time adaptation without requiring model updates, termed ”model-optimization-free” techniques, are increasingly garnering interest. These approaches enhance the utilization of test data by either memorizing it or incorporating additional modules on top of the original model.
Both ETLT  (Fan et al., 2022) and GOODAT  (Wang et al., 2024a) retain the integrity of the original model by training an add-on module to adjust the OOD score, rather than altering the original model itself. ETLT observes a linear correlation between the feature representation and the OOD score of a test input. In other words, for a given image, the pair consisting of its feature and OOD score (feature, OOD score) exhibits a linear correlation. Furthermore, the aforementioned pairs of ID and OOD data is linearly separable. Based on the observation, ETLT proposes to learn a linear regression module trained from the (feature, OOD score) pair. The availability of complete test datasets may not always be feasible, therefore,  Fan et al. (2022) also provide an online variant to ensure safer deployment. Similarly, GOODAT develops an add-on named graph masker, designed specifically for graph data. It integrates GIB-boosted losses and employs it as the metric for OOD scoring. In contrast, AdaOOD  (Zhang et al., 2023b) avoids any additional training burden through a non-parametric k-nearest neighbours approach. Similarly, OODD (Yang et al., 2025) constructs a dynamic OOD dictionary, managed by a priority queue, which continuously updates with feature representations of test samples based on their OOD scores. The core principle behind AdaOOD and OODD is the maintenance of a memory bank, similar to the approach taken by AUTO.
@@
Online v.s. Offline. Most post-hoc methods  (Liang et al., 2018b; Fan et al., 2022) have traditionally emphasized the offline scenario, where OOD detectors remain static and fixed after deployment. In contrast, the majority of test-time methods (Zhang et al., 2023b; Gao et al., 2023a) adopt the online scenario to obtain the decision boundary dynamically, minimizing the risk of incorrect OOD predictions at each time step.

### v4

@@ 3.2. OOD Detection Approaches with Both ID and OOD Data
Outlier Mining. The traditional OE concept assumes the existence of ID input $\mathcal{D}_{\text{in}}$ and OOD input $\mathcal{D}_{\text{out}}$, both independently and heterogeneously distributed, originating from different sources. However, this premise cannot be fully guaranteed in the current training process due to potential noise in the training OOD data. Outlier Mining differs slightly from the traditional OE approach in that, although it also utilizes real-world OOD samples to address the issue, it focuses on identifying the optimal selection within the existing OOD data. The main process is depicted in Fig.  3 (d). Different approaches have been proposed for selecting representative OOD samples: POEM  (Ming et al., 2022a) utilizes posterior sampling to identify high boundary score anomalies, while  Li and Vasconcelos (2020) employs data resampling with priority score reweighting for hard negative instances. In contrast, DAOL  (Wang et al., 2023a) models OOD distribution using Wasserstein balls to select challenging OOD samples based on the disparity between real and auxiliary data.
@@
Long-tail ID data. Recent studies have explored OOD detection in scenarios with imbalanced ID training data. Different approaches have been proposed to address this challenge: PASCAL  (Wang et al., 2022b) uses contrastive loss to separate tail-class from OOD data, while COCL  (Miao et al., 2024a) introduces a learnable tail class prototype to better distinguish between them.  Choi et al. (2023) focuses on balancing cross-class distribution of auxiliary OOD data through energy regularization, and EAT  (Wei et al., 2024) expands the classification space by dynamically assigning virtual labels to OOD data during training.
@@ 4.1. Post-hoc OOD Detection Approaches
Overview. Given a well-trained model, this problem scenario involves utilizing only the intermediate results computed by the trained model during testing, without modifying any parameters of the model, to accomplish the OOD detection task. Post-hoc methods are favored for their lightweight nature, low computational costs, and the fact that they require minimal modifications to the model and objectives. Its main objective is to construct an effective scoring function that can accurately reflect the behavior of ID data. These characteristics make them highly desirable for convenient and straightforward deployment in practical scenarios. Post-hoc approaches are categorized into five types: Output-based, Distance-based, Gradient-based, Feature-based and Density-based. Recent work on this type of problem has some recent progress. A summary of the key factors involved in such methods is given in Table  2.
@@
Output-based. Algorithms based on output primarily aim to explore the latent representations of the output from the intermediate layers of neural networks, which include logits and class distributions, among others. MSP  (Hendrycks and Gimpel, 2017) is the first to employ the maximum softmax value to validate OOD detection effectiveness. For OOD samples, their output probability distribution tends to be closer to a uniform distribution, demonstrating the model’s inability to correctly classify the category. Departing from the MSP method, MaxLogits (Hendrycks et al., 2022) detects OOD samples using the maximum logit value, while Energy (Liu et al., 2020a) employs an energy-based score related to input probability density. Compared to MaxLogits, Energy is less sensitive to model overconfidence and provides a more theoretically grounded metric.
@@
Gradient-based. Gradient-based methods leverage model gradients to quantify uncertainty. Grad (Lee and AlRegib, 2020) and GradNorm (Huang et al., 2021) both utilize gradient magnitudes, with the latter specifically comparing KL-divergence-based gradients between ID and OOD samples. GradOrth (Behpour et al., 2023), instead of focusing on overall gradient norms, projects gradients onto low-rank subspaces to capture OOD features. GAIA (Chen et al., 2023) introduces channel-wise and zero-deflation abnormality scores to assess distribution shifts without prior knowledge. OPNP (Chen et al., 2024a) further improves OOD detection by pruning parameters and neurons near zero, thus enhancing generalization. In recent research, PRO (Chen et al., 2025a) introduces an adversarial score, with the core idea being to perturb the input data using gradient descent to search for the local minimum score near the original input. This method effectively lowers the confidence of OOD samples and is compatible with existing MSP detection methods during score calculation, thereby enhancing detection performance. S&I (Zhang et al., 2025a) identifies the issue of insufficient sample gradient feature discrimination in this type of method. By introducing adversarial samples and integrating gradients along the attribution path, it achieves a more accurate feature distinction strategy.
Feature-based. Research in this direction explores the role of intermediate neural features for OOD detection. ODIN (Liang et al., 2018a), inspired by adversarial examples, demonstrates that small input perturbations can enhance OOD detection performance. The perturbation formula is as follows:
@@
where the parameter $\varepsilon$ is the perturbation magnitude. For a given input $x$, compute its logit output $p_{c}(x)$. Therefore, ReAct  (Sun et al., 2021) focuses on the high activation values in the intermediate results of the model. These activation values do not affect the model’s classification, but truncating high activation values can significantly improve OOD detection performance. VRA  (Xu et al., 2023), an extended iteration of ReAct  (Sun et al., 2021), builds upon the premise of ReAct which truncates only high activation values. However, VRA posits that this might not be the optimal solution and, therefore, employs a variational approach to seek an optimal solution. It utilizes piecewise functions to emulate suppression or amplification operations, aiding the model in recognizing anomalous data. KANs (Canevaro et al., 2025) also focus on the neuron activation states. Due to the local neuroplasticity characteristic of Kolmogorov-Arnold Networks (Liu et al., 2025b), KANs distinguish between ID and OOD data by comparing the activation response differences between the trained and untrained KANs. SHE (Zhang et al., 2022) introduces a storage-based approach by leveraging Hopfield energy (Krotov and Hopfield, 2016) to facilitate OOD detection, averaging logits per class to form a reference pattern. Building on channel selection, DDCS (Yuan et al., 2024) evaluates and corrects neural network channels according to their discriminative ability, using inter-class similarity and variance. Similarly, LINe (Yong Hyun Ahn, 2023) focuses on neuron-level feature outputs, employing Shapley value pruning to retain only the most informative neurons for OOD detection. CADRef (Ling et al., 2025) decouples the features by using a strategy based on the sign alignment between the relative features and model weights, dividing the sample features into positive and negative errors. Intuitively, the positive and negative error components influence the final output logits, which helps in identifying OOD images from this perspective. ITP (Xu and Yang, 2025) optimizes the model by quantifying the contribution of model parameters to ID data prediction, specifically by evaluating the contribution of each parameter through its partial derivatives and removing those with low contribution. Additionally, during the testing phase, ITP uses a right-tail Z-score test to assess whether any parameter exhibits overconfidence in the classification of a given test sample, further refining the model’s decision-making.
Feature shaping refines intermediate features during forward propagation, offering a simple and effective way to enhance OOD detection without affecting original classification results. Building on this, ViM (Wang et al., 2022c) integrates features, logits, and probabilities by constructing virtual logits, and leverages the penultimate layer’s null space—irrelevant to classification yet highly effective for OOD detection. Its computation of score $S$ can be expressed in the following form:
@@
where $\alpha$ is a scaling constant, computed by the model. Here $z=z^{P}+z^{P^{\perp}}$ and $z^{P^{\perp}}$ is the projection of $z$ to $P^{\perp}$. And it have $Wz^{P^{\perp}}=0$. Additionally, $LogSumExp$ represents the computation process of the energy function  (Liu et al., 2020a), and $f(z)$ represents the logit output of the model. The first term here represents virtual logits, while the second term represents the score of the energy function. Neco  (Ammar et al., 2024) subsequently reveals the prevalent phenomenon of neural collapse in contemporary neural networks, impacting OOD detection performance. The observation of orthogonal trends between ID data and OOD data features is leveraged to differentiate OOD data. Similarly, NCI (Liu and Qin, 2023) studies the phenomenon of neural collapse and discovers that during training, ID samples tend to cluster around class weight vectors in the penultimate feature space, while OOD samples do not exhibit this clustering behavior. Therefore, NCI leverages the proximity between features and class weight vectors to effectively distinguish ID samples from OOD samples, while maintaining low computational cost. ASH (Djurisic et al., 2023) and neuron activation pruning focus on direct modification or removal of activations, with ASH offering a simple dynamic scheme and pruning methods building upon its foundation. In contrast, NAC (Liu et al., 2024b) introduces neuron activation coverage as a statistical indicator, aiming to capture OOD likelihood based on rarely activated neurons. BLOOD (Jelenić et al., 2024) and related methods leverage the smoother intermediate representations of ID data, using this property to design new statistical measures for OOD discrimination. Feature shaping approaches (Zhao et al., 2024) partition the feature space and estimate piecewise approximations to logits, while SCALE (Xu et al., 2024a) highlights the importance of scaling metrics and observes lower pruning rates for OOD samples. Overall, these methods differ in whether they reshape activations directly, exploit statistical properties, or design new metrics for improved OOD detection.
Density-based. Recent density-based OOD detection methods have achieved notable improvements by more effectively modeling the true data distribution. For example, GEM (Morteza and Li, 2022) employs class-conditional Gaussian distributions with statistical metrics for model validation. Building upon this, ConjNorm (Peng et al., 2024) extends the framework to exponential family distributions using Bregman divergence, enabling broader applications.
@@ 4.2. Test-Time Adaptive OOD Detection Approaches
Overview. Recently, many works (Yu et al., 2024b, 2023; Sun et al., ; Yu et al., 2025; Liang et al., ) have focused on test-time adaptive methods, where a model, trained on the training set, adapts during the testing phase. In the context of OOD detection, these methods aim to leverage unlabeled test data—whether the complete test set or mini-batches—to improve performance through model adaptation. Test-time adaptive approaches are based on the theoretical insight  (Fang et al., 2022) that detecting OOD samples using only ID samples without any additional knowledge can be challenging and potentially limited. These methods can be divided into two categories according to whether the model be modified during the testing time: Model-optimization-based and Model-optimization-free. Both of them undergo a post-training phase, during which the trained model can be adapted, regardless of whether it is updated.
@@
A line of methods (Katz-Samuels et al., 2022; Du et al., 2024a) leverage “wild data”—a mix of unlabeled ID and OOD samples—to enhance OOD detection. WOODS (Katz-Samuels et al., 2022) focuses on cleaning wild data for reliable OOD regularization, while SAL (Du et al., 2024a) analyzes its impact through separability and learnability. However, wild data may introduce unintended information if it differs from test OOD, and additional training requirements can increase computational costs.
Another line of model-optimization-based methods draw inspiration from semi-supervised-learning techniques  (Zhu, 2005) for efficient post-training. While basic pseudo-labeling  (Arazo et al., 2020) labels test data directly, AUTO  (Yang et al., 2023a) uniquely focuses on pseudo-OOD data with semantic consistency. In contrast, ATTA  (Gao et al., 2023a) and SODA  (Geng et al., 2023) utilize both pseudo-ID and pseudo-OOD data, with SODA implementing dual-loss optimization and ATTA applying differential weighting strategies.
@@
Model-optimization-free. Modifying the original trained model is infeasible in certain security-sensitive scenarios. Therefore, methods enabling test-time adaptation without requiring model updates, termed “model-optimization-free” techniques, are increasingly garnering interest. These approaches enhance the utilization of test data by either memorizing it or incorporating additional modules on top of the original model. Both ETLT  (Fan et al., 2022) and GOODAT  (Wang et al., 2024a) preserve model integrity by training add-on modules for OOD score adjustment. ETLT leverages the linear correlation between features and OOD scores, implementing both offline and online variants, while GOODAT develops a graph-specific masker with GIB-boosted losses. In contrast, AdaOOD  (Zhang et al., 2023b) and OODD  (Yang et al., 2025) adopt non-parametric approaches, utilizing memory banks with k-nearest neighbours and dynamic dictionaries, respectively.
@@
Online v.s. Offline. Most post-hoc methods  (Liang et al., 2018b; Fan et al., 2022) emphasize the offline scenario, where OOD detectors remain static and fixed after deployment. In contrast, the majority of test-time methods (Zhang et al., 2023b; Gao et al., 2023a) adopt the online scenario to obtain the decision boundary dynamically, minimizing the risk of incorrect OOD predictions at each time step.